A rectangle has a width of 15 feet and a diagonal of a length 22 feet. How long is the rectangle? What is the perimeter of the rectangle? I know the formulas to determine the perimeter, the area, and the diagonal. I am just confused of where to start. Greatly appreciate the assistance. I think I can just get started I would be okay. By the way, this not a textbook question.
Answer by stanbon(75887)

A rectangle has a width of 15 feet and a diagonal of a length 22 feet. How long is the rectangle? What is the perimeter of the rectangle? I know the formulas to determine the perimeter, the area, and the diagonal.
-------------------
Draw the rectangle with its diagonal; label the width 15 and the diagonal 22.
Use Pythagoras to solve for the length:
22^2 = 15^2 + length^2
length^2 = 22^2-15^2 = 259
Take the square root of both sides to get:
length = 16.09 ft
===================
Cheers,
Stan H.